In a geometric sequence, the common ratio is -5. The sum of the first 3 terms is 147.What is the value of the first term in the sequence?
Answer by stanbon(75887)

In a geometric sequence, the common ratio is -5. The sum of the first 3 terms is 147.What is the value of the first term in the sequence?
--------
a + ar + ar^2 = 147
a(1+r+r^2) = 147
a(1-5+25) = 147
a = 147/21
a = 7
-------------
Cheers,
Stan H.
---------